OpenStudy (anonymous):

what are the solutions for x^2+37=2x

OpenStudy (anonymous):

\[x^2 - 2x +37 = 0\]\[Quadratic Formula: (-b \pm \sqrt{b^2-4ac})/2a\] a= 1 b = -2 c=37 This does not have real solutions :D